2. The controller of Hall Industries has collected the following monthly expense data for use in analyzing the cost behavior of maintenance costs.

total total
Month Maintenance Costs Machine Hours
1 2460 3500
2 3000 4000
3 3600 6000
4 4500 7900
5 3200 5000
6 4620 8000

(a) Determine the fixed- and variable-cost components using the high-low method.

.................................500 P18.1A (LO1, 2) The controller of Rather Production has collected the following monthly expense data for analyzing the co
Answer #1

Solution 2:

Variable cost per machine hour using high low method = (Cost at high level - Cost at low level) / (High level activity - Low level activity)

= ($4,620 - $2,460) / (8000 - 3500)

= $0.48 per machine hour

Fixed costs = $4,620 - (8000*$0.48) = $780

Solution 18a:

Variable cost per machine hour using high low method = (Cost at high level - Cost at low level) / (High level activity - Low level activity)

= ($5,860 - $2,500) / (720 - 300)

= $8 per machine hour

Fixed costs = $5,860 - (720*$8) = $100

Solution 18b:

Estimated cost at 500 machine hours = $100 + 500*$8 = $4,100

Amount differ from march cost = $4,100 - $3,600 = $500